OpenStudy (anonymous):

Evaluate the rational function or expression for x=1 and x=-2. write undefined if appropriate. y=1/x-5

OpenStudy (anonymous):

y = 1/(x - 5) for x=1: y = 1/(1 - 5) = 1/-4 = -1/4 for x=-2: y = 1(-2 - 5) = 1/-7 = -1/7